Ce nter City East Parking Garage has a capacity of 251 cars more than Center City West Parking Garage. If the combined capacity for the two garages is 1221 cars, find the capacity for each garage
Answer by Maths68(1474)

Ce nter City East Parking Garage has a capacity of 251 cars more than Center City West Parking Garage. If the combined capacity for the two garages is 1221 cars, find the capacity for each garage
Let
The Capacity of Center City West Parking = x
The Capacity of Center City East Parking = x + 251
Parking Capacity of Both Centres = Total Parking Capacity
x+x+251=1221
2x+251=1221
2x=1221-251
2x=970
2x/2=970/2
x=485


The Capacity of Center City West Parking = x = 485
The Capacity of Center City East Parking = x + 251 = 485+251=736


Check
=====
Parking Capacity of Both Centres = Total Parking Capacity
485+736=1221
1221=1221